Circular No. 7645

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cp IN UGC 10738 M. Modjaz and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on June 19.3 (mag ∼ 16.5) and 20.3 UT (mag ∼ 16.3). SN 2001cp is located at α = 17^{h}11^{m}02^{s}.58, δ = +5^{o}50′26′′.8 (equinox 2000.0), which is 49′′.1 west and 40′′.0 south of the nucleus of UGC 10738. A KAIT image taken on June 12.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001ck IN UGC 9425 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001ck (cf. IAUC 7441), obtained by P. Berlind on June 17.28 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum light. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 10 408 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 3 ± 2 days after maximum light.

SUPERNOVA 2001cm IN NGC 5965 Matheson et al. add that a spectrum of SN 2001cm (cf. IAUC 7441),

obtained as above on June 17.30 UT, shows it to be a type-II supernova near maximum light. The spectrum consists of a very blue continuum with superposed P-Cyg lines of hydrogen. Adopting the NED recession velocity of 3412 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 10 000 km/s.

COMET C/2001 A2 (LINEAR) Visual m_{1} and coma-diameter estimates: May 11.44 UT, 5.3, 7′ (A.

Pearce, Nedlands, W. Australia, 20×80 binoculars); 21.40, 5.1, 7′ (M. Mattiazzo, Wallaroo, S. Australia, 7×50 binoculars); June 5.37, 4.9, 12′ (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars); 6.90, 4.6, 7′ (Pearce); 11.91, 3.6, 14′ (Pearce, 8×40 binoc.); 12.82, 3.3, 12′ (Mattiazzo, naked eye; moonlight); 15.33, 3.5, 15′ (W. Souza, Joaquim Egidio, Brazil, 8×30 binoculars); 17.32, 3.9, – (Aguiar, naked eye); 18.79, 4.3, 15′ (Mattiazzo; with 25×100 binoculars, tail > 4^{o}.5 long in p.a. 219^{o}); 19.90, 4.6, 14′ (Pearce).

2001 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green